# Consistency of JWST Black Hole Observations with NANOGrav Gravitational Wave Measurements

### v1

@@ Abstract
Effects that potentially determine the dynamics of the binary at the intermediate scales include gas infall and viscous drag. The timescale of the binary evolution by gas infall can be estimated from the accretion timescale Begelman et al. (1980), $t_{\rm gas}\simeq m_{1}/\dot{m}_{1}$, where $m_{1}$ is the mass of the heavier BH, and the timescale of viscous drag has been estimated in Ivanov et al. (1999) to scale as, $t_{\rm vd}\propto\sqrt{d}/M^{1/14}$. These correspond, respectively, to $(\alpha,\beta)=(8/3,5/8)$ and $(\alpha,\beta)\approx(7/3,3/4)$, while a transition directly from the loss-cone scattering to GW-driven evolution would correspond to $(\alpha,\beta)=(10/3,2/5)$. For the PTA fit shown in Fig. 3 we fixed $\beta=5/8$ after checking that the fit is not very sensitive to the value of $\beta$. We also find that the above values of $\alpha$ provide good fits to the data.
@@
Figure 5: The detectability probability $p_{\rm det}$ (left) arising from a threshold on the flux and the dual identification probability $p_{\rm dual}$ (right) arising from a threshold on the angular separation.
Given the estimates of the effective timescales, the expected number of dual AGNs that JWST should detect is
$N_{\rm AGN,2}=f_{\rm sky}\int\!{\rm d}\lambda\frac{{\rm d}d}{|\dot{d}|}\,p_{\rm det}(m_{1},z)p_{\rm det}(m_{2},z)p_{\rm dual}(d,z)\,,$ (12)
where $f_{\rm sky}$ is the sky fraction being observed, $p_{\rm det}$ is the probability that the BH is luminous enough to be detectable, $p_{\rm dual}$ is the probability that a dual AGN is not misinterpreted as a single AGN and
${\rm d}\lambda\equiv\frac{1}{1+z}\frac{{\rm d}V_{c}}{{\rm d}z}\frac{{\rm d}R_{\rm BH}}{{\rm d}m_{1}{\rm d}m_{2}}{\rm d}m_{1}{\rm d}m_{2}{\rm d}z\,.$ (13)
We assume that the AGN activity is triggered when the separation between the centers of the galaxies is $d<0.1\min(R_{{\rm v},1},R_{{\rm v},2})$ where $R_{{\rm v},j}$ denote the virial radii of the galaxies. We estimate the probability $p_{\rm det}$ by setting a threshold on the flux
@@
where $D_{L}$ denotes the luminosity distance of the system. We approximate the AGN luminosity optimistically as $L\simeq L_{\rm Edd}\approx 1.3\times 10^{44}\,(m_{\rm BH}/10^{6}M_{\odot})\,{\rm erg}/{\rm s}$ and choose $F_{\rm sens}\simeq 10^{-15}\,{\rm erg}/{\rm s}/{\rm cm}^{2}$, which corresponds to the dimmest source found in Perna et al. (2023). The threshold on the flux gives a redshift dependent lower bound on the BH mass, as shown in the left panel of Fig. 5. Moreover, based on Perna et al. (2023), we consider $z>3$.
@@
The probability $p_{\rm dual}$ is shown in the right panel of Fig. 5. The AGNs for which the accompanying SMBH is not luminous enough and those luminous dual AGNs for which the angular separation is too small to be resolved are seen as single AGNs, whose expected number is given by
$N_{\rm AGN,1}$ $=f_{\rm sky}\!\int{\rm d}\lambda\frac{{\rm d}d}{|\dot{d}|}\bigg{[}p_{\rm det}(m_{1},z)(1-p_{\rm det}(m_{2},z))$ (17)
$+(1-p_{\rm det}(m_{1},z))p_{\rm det}(m_{2},z)$
$+p_{\rm det}(m_{1},z)p_{\rm det}(m_{2},z)(1-p_{\rm dual}(d,z))\bigg{]}\,.$
The GA-NIFS survey has reported a sample of $\bar{N}_{1}=12$ single AGNs and $\bar{N}_{2}=4$ dual AGNs Perna et al. (2023). We write the timescales of dynamical friction and stellar loss-cone scattering as $t_{\rm dyn}=20\,{\rm Myr}\,A_{\rm dyn}[M/10^{9}M_{\odot}]^{-1}[d/{\rm kpc}]^{2}$ and $t_{\rm sh}=300\,{\rm Myr}\,A_{\rm sh}[d/{\rm pc}]^{-1}$ and fit the parameters $\boldsymbol{\theta}=(\alpha,f_{\rm ref},A_{\rm dyn},A_{\rm sh})$ using the likelihood
$\mathcal{L}(\boldsymbol{\theta})=P(N_{\rm AGN,1}/N_{\rm AGN,2}(\boldsymbol{\theta}))\,,$ (18)
where
$P(x)=\int{\rm d}N_{1}{\rm d}N_{2}\,P(a|\bar{N}_{1})P(b|\bar{N}_{2})\delta(x-N_{2}/N_{1})$ (19)
is the probability density of the fraction of two Poisson distributed random variables.
The upper panel of Fig. 6 illustrates the more important environmental energy loss mechanisms for a binary system of SMBHs with masses $M=10^{9}M_{\odot}$ as functions of their separation, and displays the separation ranges that are probed by PTAs and by JWST observations Perna et al. (2023) of dual AGNs. The lower panel of Fig. 6 shows posteriors for the parameters $(A_{\rm dyn},A_{\rm sh})$ that characterize the timescales of dynamical friction and stellar loss-cone scattering, as discussed in the text, which yield a dual AGN fraction compatible with the JWST data Perna et al. (2023). The values for $\alpha$ and $f_{\rm ref}$ are the best fit values to the NANOgrav data, which illustrates that is possible to get a coherent evolution of the binary that explain both the spectral shape of NANOgrav as well as the high abundance of dual AGNs reported by JWST.
Figure 6: Upper panel: Illustration of the more important environmental energy loss mechanisms for a binary system of SMBHs with masses $m_{\rm BH}=10^{9}M_{\odot}$ as functions of their separation. The vertical bands illustrate the separation ranges that are probed by PTAs and by JWST observations Perna et al. (2023) of single and dual AGNs. Lower panel: Posteriors for the parameters $(A_{\rm dyn},A_{\rm sh})$ that characterize the timescales of dynamical friction and stellar loss-cone scattering, as discussed in the text, for a binary system of SMBHs with masses $M=10^{9}M_{\odot}$ and the best fit parameters for the environmental effects to fit NANOgrav background $\alpha=2$ and $f_{\rm ref}=15$ nHz.
Conclusions – We have presented in this paper a global analysis of the available data on high-$z$ SMBHs detected with the JWST Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. ; Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023) and in previous observations Izumi et al. (2021) in conjunction with results from previous samples of SMBHs in low-$z$ AGNs and inactive galaxies Reines and Volonteri (2015) and measurements of the SGWB background by NANOGrav Agazie et al. (2023a), which are compatible with data from other PTAs Agazie et al. (2023c).
@@
The JWST data are compatible with previous high-$z$ SMBH data, and both are compatible with a sample of low-$z$ inactive galaxies (IGs), but are quite incompatible with a low-$z$ AGN sample – see Fig. 2 and Table 1. Moreover, we have shown that our global fit to the high-$z$ SMBH and local IG data also fits well the NANOGrav data on the nHz SGWB, whereas alternative interpretations of the high-$z$ SMBH data that invoke evolution with $z$ of the $M_{\rm BH}/M_{\rm*}$ relationship Pacucci and Loeb (2024) or selection biases and measurement uncertainties Li et al. (2024) provide only poor fits to the SGWB data, as seen in Fig. 4.
@@
We have also shown how the dual AGN fraction reported in Perna et al. (2023) can be accommodated within our global analysis. Our interpretation of the data on the dual AGN fraction Perna et al. (2023) is related to the need for environmental effects to accommodate the NANOGrav data, see Fig. 3 and the upper panel of Fig. 4 as well as Agazie et al. (2023b) and Ellis et al. (2024a). Fig. 6 shows that an extrapolation of these environmental effects to larger scales can accommodate the fraction of dual AGNs detected with JWST.
@@
Observations of high-$z$ SMBHs are still in their infancy, with much more data to come from JWST and other sources. Likewise, PTA data on the nHz SGWB will develop rapidly in the near future. Doubtless both sets of data will provide new puzzles. The central point of our paper is that these data sets should be considered together, and that their puzzles may share common features. Resolution of these puzzles may also require inputs from other observational programmes such as measurements of GWs at higher frequencies, interpreted within an integrated approach.
@@ Acknowledgements.
Acknowledgments – This work was supported by the European Regional Development Fund through the CoE program grant TK133 and by the Estonian Research Council grants PSG869, PRG803 and RVTT7. The work of J.E. and M.F. was supported by the United Kingdom STFC Grants ST/T000759/1 and ST/T00679X/1. The work of V.V. was partially supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101065736.
@@ Appendix A Appendix: Compilation of JWST Data and Corner Plots
Figure 7: Scatter plot in the $(m_{\rm BH}/M_{\odot},M_{*}/M_{\odot})$ plane of the JWST data analyzed in this paper. The following are the papers listed in the legend: Übler+ Übler et al. (2023), Larson+ Larson et al. (2023), Harikane+ Harikane et al. (2023), Bogdan+ Bogdan et al. , Ding+ Ding et al. (2023), Maiolino+ Maiolino et al. (2023), Yue+ Yue et al. (2023).

### v2

@@ Abstract
Effects that potentially determine the dynamics of binaries at intermediate scales include gas infall and viscous drag. The timescale of the binary evolution by gas infall can be estimated from the accretion timescale Begelman et al. (1980), $t_{\rm gas}\simeq m_{1}/\dot{m}_{1}$, where $m_{1}$ is the mass of the heavier BH, and the timescale of viscous drag has been estimated in Ivanov et al. (1999) to scale as, $t_{\rm vd}\propto\sqrt{d}/M^{1/14}$. These correspond, respectively, to $(\alpha,\beta)=(8/3,5/8)$ and $(\alpha,\beta)\approx(7/3,3/4)$, while a transition directly from the loss-cone scattering to GW-driven evolution would correspond to $(\alpha,\beta)=(10/3,2/5)$. For the PTA fit shown in Fig. 3 we fixed $\beta=5/8$ after checking that the fit is not very sensitive to the value of $\beta$. We also find that the above values of $\alpha$ provide good fits to the data.
Given the estimates for the effective timescales, we can estimate the expected number of dual and single AGNs detected by JWST. We assume that AGN activity is triggered only during major mergers where the stellar mass of the heavier galaxy is not more than three times the stellar mass of the lighter one. In minor mergers only the SMBH in the lighter galaxy typically becomes active and its activity ceases quickly as the larger galaxy rips apart the smaller one. This is in agreement with numerical simulations Scudder et al. (2012); Callegari et al. (2011); Capelo et al. (2015) and low-$z$ observations Treister et al. (2012); Comerford et al. (2015); Donley et al. (2018); Koss et al. (2012), except for a small number of objects Liu et al. (2018); Mićić et al. (2023). Dual AGNs in which one of the AGNs is not luminous enough or whose angular separation is not large enough to be resolved and AGNs generated in mergers where the other galaxy did not contain an SMBH are seen as single AGNs.
The expected number of detectable dual AGNs is
$\bar{N}_{\rm AGN,2}=f_{\rm sky}\int\!{\rm d}\lambda_{\rm 2}\,p_{\rm det}(m_{1},z)p_{\rm det}(m_{2},z)p_{\rm dual}(d,z)\,,$ (12)
where $f_{\rm sky}$ is the sky fraction being observed, $p_{\rm det}$ the probability that the BH is luminous enough to be detectable, and $p_{\rm dual}$ is the probability that a dual AGN is resolvable. Counting all mergers where at least one of the galaxies is occupied by an SMBH that becomes active during the merger and is sufficiently luminous to be detected, the expected number of detectable AGNs is
$\bar{N}_{\rm AGN,tot}=$   $\,\,2f_{\rm sky}\int{\rm d}\lambda_{1}\,p_{\rm det}(m,z)$ (13)
$+2f_{\rm sky}\int{\rm d}\lambda_{2}\,p_{\rm det}(m_{2},z)$
$-f_{\rm sky}\int{\rm d}\lambda_{2}\,p_{\rm det}(m_{1},z)p_{\rm det}(m_{2},z)\,.$
The merger rate and the evolution of the BH separation enter through
${\rm d}\lambda_{1}\equiv\frac{1}{1+z}\frac{{\rm d}V_{c}}{{\rm d}z}\frac{{\rm d}n_{\rm AGN}^{1}}{{\rm d}m{\rm d}d}{\rm d}m{\rm d}z{\rm d}d$ (14)
and
${\rm d}\lambda_{2}\equiv\frac{1}{1+z}\frac{{\rm d}V_{c}}{{\rm d}z}\frac{{\rm d}n_{\rm AGN}^{2}}{{\rm d}m_{1}{\rm d}m_{2}{\rm d}d}{\rm d}m_{1}{\rm d}m_{2}{\rm d}z{\rm d}d\,,$ (15)
where
$\frac{{\rm d}n_{\rm AGN}^{1}}{{\rm d}m{\rm d}d}=\frac{1-\sqrt{p_{\rm BH}}}{|\dot{d}|}\int\!{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}\frac{{\rm d}R_{h}}{{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}}$ (16)
×$\times p_{\rm AGN}(M_{{\rm v},1},M_{{\rm v},2},d)\frac{{\rm d}p_{\rm occ}(m|M_{*}(M_{{\rm v},2},z))}{{\rm d}m}$
is the number density of AGNs formed in galaxy mergers where one of the galaxies did not contain an SMBH, and
$\frac{{\rm d}n_{\rm AGN}^{2}}{{\rm d}m_{1}{\rm d}m_{2}{\rm d}d}=\frac{1}{|\dot{d}|}\int\!{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}\frac{{\rm d}R_{h}}{{\rm d}M_{{\rm v},1}{\rm d}M_{{\rm v},2}}$ (17)
×$\!\times\!p_{\rm AGN}(M_{{\rm v},1},M_{{\rm v},2},d)\!\!\prod_{j=1,2}\!\frac{{\rm d}p_{\rm occ}(m_{j}|M_{*}(M_{{\rm v},j},z))}{{\rm d}m_{j}}$
is the number density of dual AGNs. The probability that a galaxy with stellar mass $M_{*}$ is occupied by an SMBH with mass $m$ can be expressed as
$\frac{{\rm d}p_{\rm occ}(m|M_{*})}{{\rm d}m}=\sqrt{p_{\rm BH}}\,\frac{{\rm d}P(m|M_{*})}{{\rm d}m}$ (18)
where the distribution ${\rm d}P/{\rm d}m$, given by Eq. (1), is normalized to 1, so $\sqrt{p_{\rm BH}}$ can be interpreted as the SMBH occupation fraction, which we take to be independent of $M_{*}$. The prefactor $1-\sqrt{p_{\rm BH}}$ in (LABEL:eq:nAGN1) gives the probability that one of the merging galaxies does not contain an SMBH. In this case, dynamical friction drives this single SMBH towards the centre of the galaxy that resulted from the merger. So, for Eq. (LABEL:eq:nAGN1) we estimate the timescale as $t_{\rm eff}=t_{\rm dyn}$. We assume that the AGN activity starts when the distance between the galaxies becomes smaller than the half-stellar radius of the smaller galaxy. We approximate the half-stellar radius by $10\%$ of the virial radius of the halo:
$p_{\rm AGN}(M_{{\rm v},1},M_{{\rm v},2},d)=\Theta(0.1\min_{j}R_{{\rm v},j}-d)$ (19)
×$\hskip 40.0pt\times\Theta(3\min_{j}M_{*,j}-\max_{j}M_{*,j})\,.$
In the right panel of Fig. 5 the shades of orange show the dual AGN probability ${\rm d}n_{\rm AGN}^{2}/{\rm d}n_{\rm tot}$ where ${\rm d}n_{\rm tot}$ is obtained from Eq. (LABEL:eq:nAGN2) with $p_{\rm AGN}=1$.
@@
Figure 5: Left panel: The detectability probability $p_{\rm det}$ arising from the JWST sensitivity threshold on the flux. Right panel: The red regions show the dual identification probability $p_{\rm dual}$ at $z=3$ arising from the JWST angular resolution. The orange regions show the probability that both galactic nuclei are active for $m_{\rm BH,2}=10^{8}M_{\odot}$ at $z=3$.
We estimate the detection probability $p_{\rm det}$ by setting a threshold on the flux
@@
where $D_{L}$ denotes the luminosity distance of the system. We approximate the AGN luminosity optimistically as $L\simeq L_{\rm Edd}/3\approx 4.3\times 10^{43}\,(m_{\rm BH}/10^{6}M_{\odot})\,{\rm erg}/{\rm s}$ and choose $F_{\rm sens}\simeq 10^{-15}\,{\rm erg}/{\rm s}/{\rm cm}^{2}$, which corresponds to the dimmest source found in Perna et al. (2023). The threshold on the flux gives a redshift-dependent lower bound on the BH mass, as shown in the left panel of Fig. 5.
@@
The probability $p_{\rm dual}$ is shown in the right panel of Fig. 5.
Figure 6: Upper panels: The expected total number of AGNs at $z>2$ and the expected dual fraction at $z>3$ as a function of the reference frequency below which the binaries are driven by GW emission, assuming $p_{\rm BH}=0.2$, $\alpha=8/3$, $A_{\rm dyn}=0.001$ and $A_{\rm sh}=1$. The dashed and solid curves are with and without galaxy mergers where one of the galaxies does not contain an SMBH. Lower panel: Illustration of the timescales of the environmental energy loss mechanisms of SMBH binaries as functions of their separations for $\alpha=8/3$, $A_{\rm dyn}=10^{-3}$ and $A_{\rm sh}=1$. The vertical bands illustrate the separation ranges probed by the PTAs and by the JWST dual AGN observations. The colour coding of the curves matches that shown in the upper panels.
[…]
In the upper panels of Fig. 6 we show the dependencies of the numbers of AGNs and fractions of dual AGNs on $f_{\rm ref}$ for $\alpha=8/3$, corresponding to the gas infall timescale, $A_{\rm sh}=1$, $A_{\rm dyn}=0.001$ and $p_{\rm BH}=0.2$. For these parameters, the lower panel of Fig. 6 illustrates the timescales of the environmental energy loss mechanisms, including GW emission, gas effects, stellar loss-cone scattering and dynamical friction, as functions of the binary separation. The different curves in the lower panel correspond to different values of $f_{\rm ref}$ indicated by the coloured dots in the upper panel. The horizontal bands in the upper panel show the 95% CL bands of the JWST LRD and dual AGN observations calculated assuming the Poisson and binomial distribution, respectively. The value of $p_{\rm BH}$ is fixed so that, neglecting AGNs that may form in mergers where only one of the galaxies contains an SMBH, the expected total AGN number roughly matches the LRD observations for large $f_{\rm ref}$. We see in the right panel that, in this benchmark case, the dual fraction is within the 95% range of the observed value for $f_{\rm ref}>6$  nHz if only the AGNs that may form in mergers where both of the galaxies contain an SMBH are considered. These values of $p_{\rm BH}$, $\alpha$ and $f_{\rm ref}$ are within the 95% CL region of the NANOGrav fit, illustrating that it is possible to obtain a coherent description of the evolution of binary systems that explains both the NANOGrav GW signal and the JWST dual AGN and LRD observations.
Fig. 6 shows that the expected number of systems seen as single AGNs would be much larger than observed if $f_{\rm ref}$ is small, because in that case, the binaries would spend a lot of time orbiting with $d\sim 1$ pc separation. This is related to the well-known final-parsec problem. Both the JWST and PTA observations prefer large values of $f_{\rm ref}$ for which the final-parsec problem is avoided. Then, the main contribution to the total number of AGNs comes from the same region as for dual AGNs, $d\gtrsim 1$ kpc, where the effective timescale is largest. In this regime, both $\bar{N}_{\rm AGN,2}$ and $\bar{N}_{\rm AGN,tot}$ are proportional to the prefactor $A_{\rm dyn}$ of the dynamical friction timescale. For the benchmark shown in Fig. 6 we have fixed $A_{\rm dyn}=0.001$ so that the result is consistent with the LRD observations for values of $p_{\rm BH}$ consistent with the NANOGrav observations.
@@
Conclusions – We have presented a global analysis of the available data on high-$z$ SMBHs detected with the JWST Übler et al. (2023); Larson et al. (2023); Harikane et al. (2023); Bogdan et al. (2024); Ding et al. (2023); Maiolino et al. (2023); Yue et al. (2023) and in previous observations Izumi et al. (2021) in conjunction with results from previous samples of SMBHs in low-$z$ AGNs and IGs Reines and Volonteri (2015) and measurements of the SGWB background by NANOGrav Agazie et al. (2023a), which are compatible with data from other PTAs Agazie et al. (2023c). The JWST data are compatible with previous high-$z$ SMBH data, and both are compatible with a sample of low-$z$ IGs, but are quite incompatible with a low-$z$ AGN sample – see Fig. 2 and Table 1. Moreover, we have shown that our global fit to the high-$z$ SMBH and local IG data also fits well the NANOGrav data on the nHz SGWB, whereas alternative interpretations of the high-$z$ SMBH data that invoke evolution with $z$ of the $M_{\rm BH}/M_{\rm*}$ relationship Pacucci and Loeb (2024) or selection biases and measurement uncertainties Li et al. (2024) provide only poor fits to the SGWB data, as seen in Fig. 4.
@@
We have also shown how the dual AGN fraction reported in Perna et al. (2023) can be accommodated within our global analysis. Our interpretation of the data on the dual AGN fraction Perna et al. (2023) is related to the need for environmental effects to accommodate the NANOGrav data, see Fig. 3 and the upper panel of Fig. 4 as well as Agazie et al. (2023b) and Ellis et al. (2024a). Fig. 6 shows that an extrapolation of these environmental effects to larger scales can accommodate the fraction of dual AGNs detected with JWST. Moreover, we also find that our global analysis is compatible with the abundance of “little red dots” in JWST observations reported recently Matthee et al. (2024); Kocevski et al. (2024).
@@
Observations of high-$z$ SMBHs are still in their infancy, with much more data to come from JWST and other sources. Likewise, PTA observations of the nHz SGWB will develop rapidly in the near future. Doubtless, both sets of data will provide new puzzles. The central point of our paper is that these data sets should be considered together and that these puzzles may share common features. Resolution of these puzzles may also require inputs from other observational programmes such as measurements of GWs at higher frequencies, interpreted within an integrated approach.
@@ Acknowledgements.
Acknowledgments – This work was supported by the European Regional Development Fund through the CoE program grant TK202 and by the Estonian Research Council grants PRG803, PSG869, RVTT3 and RVTT7. The work of J.E. and M.F. was supported by the United Kingdom STFC Grants ST/T000759/1 and ST/T00679X/1. The work of V.V. was partially supported by the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie grant agreement No. 101065736.
@@ Appendix A Appendix: Compilation of JWST Data and Corner Plots
Figure 7: Scatter plot in the $(m_{\rm BH}/M_{\odot},M_{*}/M_{\odot})$ plane of the JWST data analyzed in this paper. The following are the papers listed in the legend: Übler+ Übler et al. (2023), Larson+ Larson et al. (2023), Harikane+ Harikane et al. (2023), Bogdan+ Bogdan et al. (2024), Ding+ Ding et al. (2023), Maiolino+ Maiolino et al. (2023), Yue+ Yue et al. (2023).